OpenStudy (anonymous):

2x+3 over x-4 minus x-5 over x+2

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

( (x - 5) / (x + 4) ) - ( (x + 3) / (x - 4) ) = = ( (x - 5) (x - 4) - (x + 3) (x + 4) ) / ( (x + 4) (x - 4) ) = ( (x^2 - 4x - 5x + 20) - (x^2 + 4x + 3x + 12) ) / ( (x + 4) (x - 4) ) = (x^2 - 4x - 5x + 20 - x^2 - 4x - 3x - 12) / ( (x + 4) (x - 4) ) = (-16x + 8) / ( (x + 4) (x - 4) ) = 8 (-2x + 1) / ( (x + 4) (x - 4) ) = 8 (-2x + 1) / (x^2 - 16)

OpenStudy (anonymous):

Ill post a picture

OpenStudy (anonymous):

its number 17, A and B are just multiple choice answers

OpenStudy (anonymous):

oh, and were just subtracting those

OpenStudy (anonymous):

i know it's not the same equation, i was more so just providing an example.

OpenStudy (anonymous):

oh alright

OpenStudy (anonymous):

sry, if i misled you. i'll help you to understand the material, but i won't do the work. there's no room for learning that way.